OpenStudy (anonymous):

A new homeowner buys 120 square feet of carpet for a room in her new house. If the room is square, what is the length of each side of the room? (Recall that the area of a square is calculated by l^2 where l is the length of one side.)

OpenStudy (anonymous):

area of the square=a^2 a=the lenght of each side

OpenStudy (anonymous):

a=\[\sqrt{120}\]

OpenStudy (anonymous):

thnx